A sample of gas has a mass of 0.545 g. Its volume is 125 mL at a temperature of 95 ∘ C and a pressure of 762 mmHg. Find the molar mass of the gas.
Answer #1

mass of gas (w)=0.545g .. velume of gas CV) - 125 mL = 125 L = 0.125 L, Tood temporat use = 950 = 273+95 k = 368k Pressure =